A Survey of 118 persons was conducted at TCC, and it was found at 24 persons carried a cell phone, 59 persons carried a tablet computer and 12 carried both a cell phone and a tablet.

  1. how many people carried a cell phone or a tablet?
  2. how many people carry neither a cell phone nor a tablet?
  3. how many people carried a cell phone only?
  4. how many people carried a tablet but not a cell phone?




1
Expert's answer
2021-08-29T17:57:50-0400

Solution.

1. 24+59-12=83-12=71 people

2. 118-71=47 people

3. 24-12=12 people

4. 59-12=47 people